What is the frequency of the following?

(a) Second hand of a clock

(b) Minute hand of a clock

(c) Hour hand of a clock


1
Expert's answer
2021-06-14T17:27:04-0400

a)


"f=\frac{1}{1}=1\ Hz"

b)


"f=\frac{1}{60}=0.0167\ Hz"

c)


"f=\frac{1}{3600}=0.000278\ Hz"